OpenStudy (dmezzullo):

Solve for x: 3(x + 1) = −2(x − 1) + 6 1 4 5 25

OpenStudy (anonymous):

Distribute. 3(x + 1) = −2(x − 1) + 6 3x + 3 = -2x + 2 + 6 Combine like terms. 3x + 3 = -2x + 2 + 6 3x + 3 = -2x + 8 3x = -2x + 5 5x = 5 x = 5 /5 x = 1

OpenStudy (anonymous):

@dmezzullo

OpenStudy (dmezzullo):

tyvm

OpenStudy (dmezzullo):

medal lol

OpenStudy (anonymous):

Welcome. (:

OpenStudy (dmezzullo):

u da bomb